# Feature Engineering is Not Dead: Reviving Classical Machine Learning with Entropy, HOG, and LBP Feature Fusion for Image Classification

### v1

@@ VII-B Results on KMNIST Dataset
We performed hyperparameter tuning of an SVM with an RBF kernel via grid search using 3-fold cross-validation. The best parameters were found to be $C=200$ and $\gamma=0.001$, yielding a cross-validation accuracy of 97.55%. On the held-out test set, the tuned RBF SVM achieved a test accuracy of 94.46%, outperforming previous classical machine learning SVM approaches on KMNIST that used raw pixels or simpler handcrafted features. Table IV summarizes the comparison.
@@
TABLE IV: Comparison with existing Classical ML-based methods on KMNIST (classical ML only)
@@
Ours (PE + Corr + HOG + LBP) 780 handcrafted features 94.46%
@@ VII-C Results on EMNIST Dataset
As shown in Table V, our method outperforms existing classical SVM pipelines applied to EMNIST datasets. The results demonstrate the effectiveness of combining multiple handcrafted features for the EMNIST handwritten character recognition task. The performance is consistent with or better than previously reported classical SVM-based methods on EMNIST, which typically utilize individual feature types such as raw pixels or HOG descriptors.
TABLE V: Comparison with existing Classical ML-based methods on EMNIST (classical ML only)
@@
Ours (PE + Corr + HOG + LBP) 780 handcrafted features 93.27%
@@ VII-D Results on CIFAR-10 Coloured Dataset
To apply our handcrafted feature pipeline to CIFAR-10, we extended our method to handle RGB color channels. Each image was split into its red, green, and blue channels. Features were extracted independently from each channel using the same feature extraction pipeline (Fig. 4). Each channel contributed a set of entropy and texture features, which were concatenated to form the final feature vector. In total, this resulted in a 1281-dimensional feature vector per image. We performed hyperparameter tuning of an SVM with an RBF kernel via grid search using 3-fold cross-validation. The best parameters were found to be $C=100$ and $\gamma=0.001$, yielding a cross-validation accuracy of 61.51%. On the held-out test set, the tuned SVM achieved an accuracy of 64.81%.
TABLE VI: Comparison with existing Classical ML-based methods on CIFAR-10 (classical ML only)
@@
Ours (PE + Corr + HOG + LBP) 1281 handcrafted RGB features 64.81%
@@
As shown in Table VI, our method outperforms existing classical SVM pipelines applied to CIFAR-10. By combining entropy-based complexity measures (including patch-based permutation entropy), correlation structure, edge and texture descriptors (HOG and LBP), we capture richer information than standard approaches that rely solely on edge or intensity features.
@@
Coates et al. [28] extracted a 4 000-dimensional feature vector by learning single-layer convolutional filters via K-means and simple pooling, achieving 79.6% accuracy on CIFAR-10. However, their approach incurs a significant offline training cost and yields features that are difficult to interpret. In contrast, our permutation-entropy + HOG + LBP pipeline produces a compact, lightweight feature descriptor that can be computed in a fully feed-forward manner without any filter learning, and whose individual components (entropy, gradient histograms, binary patterns) offer clear, intuitive interpretations. This interpretability and efficiency comes with a modest trade-off in accuracy, making our method particularly well-suited to applications with limited compute or a need for transparent feature representations.
VIII Conclusion
@@
To conclude, we presented a unified handcrafted feature extraction pipeline that integrates diverse descriptors, including temporal-complexity measures such as permutation entropy, spatial correlations, edge- and texture-based features (e.g., HOG, LBP), to construct a rich and interpretable feature representation. By applying this approach to a suite of benchmark datasets—Fashion-MNIST, KMNIST, EMNIST, and CIFAR-10—we demonstrated that shallow classifiers, particularly SVM, can achieve competitive performance without relying on deep learning architectures.
Our results show that combining complementary handcrafted features significantly outperforms traditional single-feature methods in the classical machine learning regime (Fig. 6). Notably, we achieve up to 91.15% test accuracy on Fashion-MNIST and 64.81% on CIFAR-10 using only interpretable, resource-efficient techniques. While still behind state-of-the-art deep neural networks, our method narrows the gap and offers a viable alternative for constrained environments where explainability, simplicity, or limited computational resources are priorities. These findings reinforce the relevance of classical approaches when enhanced with thoughtful feature engineering, and suggest that hybrid or complementary strategies may continue to play a role alongside deep learning in future vision systems.
@@
In the future we plan to use these features as input representations for designing next-generation neural networks. Instead of feeding raw pixels into a deep model, we could build architectures that learn from more meaningful, structured inputs—like the features we extract here. This could lead to models that are faster to train, more accurate with less data, and easier to interpret. Therefore, combining the strengths of handcrafted features with the flexibility of deep learning could pave the way for next-generation computer vision systems that are both efficient and explainable.

### v2

@@ VII-B Results on KMNIST Dataset
We performed hyperparameter tuning of an SVM with an RBF kernel via grid search using 3-fold cross-validation. The best parameters were found to be $C=200$ and $\gamma=0.001$, yielding a cross-validation accuracy of 97.55%. On the held-out test set, the tuned RBF SVM achieved a test accuracy of 94.46%, outperforming previous classical machine learning SVM approaches on KMNIST that used raw pixels or simpler handcrafted features. Table VII summarizes the comparison.
@@
TABLE VII: Comparison with existing Classical ML-based methods on KMNIST (classical ML only)
@@
PE/Corr + HOG + LBP + SVM 780 handcrafted features 94.46%
@@ VII-C Results on EMNIST Dataset
As shown in Table VIII, our method outperforms existing classical SVM pipelines applied to EMNIST datasets. The results demonstrate the effectiveness of combining multiple handcrafted features for the EMNIST handwritten character recognition task. The performance is consistent with or better than previously reported classical SVM-based methods on EMNIST, which typically utilize individual feature types such as raw pixels or HOG descriptors.
TABLE VIII: Comparison with existing Classical ML-based methods on EMNIST (classical ML only)
@@
PE/Corr + HOG + LBP + SVM 780 handcrafted features 93.27%
@@ VII-D Results on CIFAR-10 Coloured Dataset
To apply our handcrafted feature pipeline to CIFAR-10, we extended our method to handle RGB color channels. Each image was split into its red, green, and blue channels. Features were extracted independently from each channel using the same feature extraction pipeline (Fig. 4). Each channel contributed a set of entropy and texture features (total 427 features), which were concatenated to form the final feature vector (yielding 1281 = 3 × 427 features). In total, this resulted in a 1281-dimensional feature vector per image, as shown in Table IX. We performed hyperparameter tuning of an SVM with an RBF kernel via grid search using 3-fold cross-validation. The best parameters were found to be $C=100$ and $\gamma=0.001$, yielding a cross-validation accuracy of 61.51%. On the held-out test set, the tuned SVM achieved an accuracy of 64.81%.
TABLE IX: Breakdown of handcrafted feature types used in CIFAR-10 (per RGB channel and total).
Feature Type (Count rationale) Per-channel Total (R+G+B)
Row-wise PE (32 rows) 32 96
Column-wise PE (32 columns) 32 96
Row Correlation (31 adjacent row pairs) 31 93
Column Correlation (31 adjacent column pairs) 31 93
Main-diagonal PE ($8{\times}8$, stride 4 $\Rightarrow 7{\times}7$) 49 147
Anti-diagonal PE (flip + $8{\times}8$, stride 4 $\Rightarrow 7{\times}7$) 49 147
Patch-wise PE ($8{\times}8$, stride 4 $\Rightarrow 7{\times}7$) 49 147
HOG ($8{\times}8$ cells, 9 bins $\Rightarrow 4{\times}4{\times}9$) 144 432
LBP ($P=8$, $R=1$ $\Rightarrow P{+}2$ bins) 10 30
Total 427 1281
TABLE X: Comparison with existing Classical ML-based methods on CIFAR-10 (classical ML only)
@@
PE/Corr + HOG + LBP + SVM 1281 handcrafted RGB features 64.81%
@@
As shown in Table X, our method outperforms existing classical SVM pipelines applied to CIFAR-10. By combining entropy-based complexity measures (including patch-based permutation entropy), correlation structure, edge and texture descriptors (HOG and LBP), we capture richer information than standard approaches that rely solely on edge or intensity features.
Color space and channel ablation on CIFAR-10. Using our original RGB configuration (entropy+HOG+LBP on all RGB channels) we obtain 64.81% accuracy. Additionaly, we evaluated channel-selective and alternative color-space variants: restricting features to “structural” channels (YCbCr $Y$ and HSV $V$) yields 60.42%, and extracting entropy across all YCbCr channels (with HOG/LBP on $Y$) yields 61.62%. These results suggest that CIFAR-10 benefits from chroma information; discarding or limiting color channels can reduce discriminative power.
@@
A different study by Coates et al. [34] extracted a 4 000-dimensional feature vector by learning single-layer convolutional filters via K-means and simple pooling, achieving 79.6% accuracy on CIFAR-10. However, their approach incurs a significant offline training cost and yields features that are difficult to interpret. In contrast, our permutation-entropy + HOG + LBP pipeline produces a compact, lightweight feature descriptor that can be computed in a fully feed-forward manner without any filter learning, and whose individual components (entropy, gradient histograms, binary patterns) offer clear, intuitive interpretations. This interpretability and efficiency comes with a modest trade-off in accuracy, making our method particularly well-suited to applications with limited compute or a need for transparent feature representations.
VIII Future Research Direction and Preliminary Results
VIII-A Neural Networks and Hand-crafted Features
To study how the proposed handcrafted descriptors interact with modern convolutional features, we consider an additional experiment on CIFAR-10 using a lightweight CNN backbone and a hybrid (deep + classical ) representation. Further, we chose the CIFAR-10 dataset to study this idea because it is a standard, moderately challenging natural image benchmark that is widely used to evaluate lightweight CNNs and allows fair comparison with both classical and deep learning methods.
As an efficient deep baseline, we adopt MobileNetV2, a compact architecture based on depthwise separable convolutions with inverted residual blocks and linear bottlenecks. The network is initialized with ImageNet-pretrained weights and adapted to CIFAR-10 by replacing the final classification layer with a 10-class head. From the original CIFAR-10 training set (50,000 images), we construct a train/validation split using a tail split of 45,000 training and 5,000 validation images, and evaluate on the standard 10,000-image test set. CIFAR-10 images are resized to $224\times 224$ pixels and normalized with ImageNet mean and variance to match the backbone’s default input conditions.
The model is fine-tuned using Adam with cross-entropy loss for 20 epochs, with a batch size of 64. We track validation accuracy and select the best checkpoint for final evaluation. This MobileNetV2 configuration achieves a test accuracy of 92.43%, with a total fine-tuning time of 23,054 s. The penultimate layer produces 1280-dimensional embeddings, which we export for all train/validation/test samples.
TABLE XI: MobileNetV2 transfer-learning configuration on CIFAR-10.
Setting Value Notes
Backbone MobileNetV2 ImageNet-pretrained
Input size $224\times 224$ Resized CIFAR-10
Train / Val / Test 45k / 5k / 10k Tail split on train
Optimizer Adam Cross-entropy loss
Epochs / Batch size 20 / 64 Fine-tuning
Embedding dimension 1280 Penultimate layer
Train time 23,054 s CPU/GPU setup as in Sec. X
Test accuracy 92.43% Best val checkpoint
On the same CIFAR-10 splits, we compute our handcrafted feature vector, resulting in a 1281-dimensional descriptor per image. For each sample, we concatenate the MobileNetV2 embedding (1280 dimensions) with the handcrafted descriptor (1281 dimensions), forming a fused feature vector of dimension 2561. This final feature vector is hybrid in nature because it combines learned deep embeddings from MobileNetV2 with explicitly defined handcrafted descriptors bringing together semantic representation and interpretable edge/texture/complexity cues in a single feature space.
Finally, a linear SVM with squared hinge loss is trained in this fused space. The hyperparameters (regularization parameter $C$, tolerance, and class weighting) are tuned by using the Optuna-based hyperparameter search in the training/validation split. The best configuration (class_weight = balanced, $C\approx 6.7\times 10^{-4}$, tol $\approx 4.9\times 10^{-5}$) yields a test accuracy of 93.10%, with an SVM training time of 313 s.
TABLE XII: Deep baseline vs. hybrid deep + handcrafted fusion on CIFAR-10.
Setting Value Notes
MobileNetV2 (fine-tuned)
Feature dimension 1,280 Deep embedding
Classifier Softmax head End-to-end training
Train time 23,054 s As measured in our setup
Test accuracy 92.43% CPU-only
MobileNetV2 + handcrafted fusion
Feature dimension 2,561 1,280 (deep) + 1,281 (handcrafted)
Classifier SVM Trained on fused features
Train time 313 s Classifier training only
Test accuracy 93.10% CPU-only
The hybrid representation improves CIFAR-10 performance from 92.43% (MobileNetV2 fine-tuning alone) to 93.10% (MobileNetV2 embeddings fused with handcrafted features and classified by SVM), corresponding to a +0.67 percentage-point gain. At this accuracy level, where reaching above 90% on CIFAR-10 is already non-trivial, such an improvement is meaningful and suggests that the proposed handcrafted descriptors provide complementary information to the deep embeddings. As a modest but meaningful gain at this accuracy level, this result is encouraging, and future work may explore similar deep + handcrafted fusion on a few other datasets, although a full-scale exploration is beyond the scope of this paper.
VIII-B Understanding efficiency
In this subsection, we provide a quantitative comparison of the computational cost of the proposed handcrafted-feature pipeline and a deep learning (CNN) baseline we saw previously. We report separate timing measurements for feature extraction and classifier training on CIFAR-10, and we analyze how these components contribute to the overall efficiency of purely classical, purely deep, and hybrid deep+handcrafted configurations.
Here, the experiments were conducted on a high-performance Linux workstation running Ubuntu 22.04.5 LTS with an Intel Xeon Platinum 8280 CPU (112 cores, 4.0 GHz) and 772,640 MiB of system memory, using CPU-only computation throughout.
Let us state the results. From a classical machine learning perspective, the SVM trained on our handcrafted features and achieving 64.81% accuracy on CIFAR-10 is competitive, because it operates on a compact, engineered 1,281-dimensional descriptor rather than raw pixels. As we show earlier in the paper, an SVM directly on pixel intensities performs noticeably worse, so the gain comes from feature engineering, not from the classifier alone. Moreover, the total cost of this purely classical pipeline, feature extraction plus SVM training, remains modest on CPU as both stages are feed-forward, parallelizable, and do not require any backpropagation through a large network.
The utility of the handcrafted descriptor becomes more evident in a hybrid setting. Once deep embeddings from a lightweight CNN (MobileNetV2) are available, concatenating these 1,280-dimensional embeddings with our 1,281-dimensional handcrafted features and training a linear SVM on the fused 2,561-dimensional space not only surpasses the CNN’s standalone test accuracy (93.10% vs. 92.43%), but also requires only a small fraction of the training time compared to end-to-end CNN fine-tuning. In this regime, the costly deep training is performed once to obtain reusable embeddings, and the handcrafted features provide an efficient, CPU-friendly way to further improve accuracy.
TABLE XIII: Computational cost comparison on CIFAR-10 (CPU-only). Feature extraction times are measured over the full dataset (train + validation + test); classifier training times are measured on the train/validation split.
Setting Value Notes
MobileNetV2 (end-to-end fine-tuning)
Feature dim. 1,280 Concatenation
Feature ext. time (s) $\approx 93$ Full dataset
Classifier train time (s) 23,054 Train/val split
Test acc. (%) 92.43% CPU-only
Handcrafted features + SVM
Feature dim. 1,281 Concatenation
Feature ext. time (s) 898 Full dataset
Classifier train time (s) 4,950 Train/val split
Test acc. (%) 64.81% CPU-only
MobileNetV2 + handcrafted + SVM
Feature dim. 2,561 Concatenation
Feature ext. time (s) $898+93$ Full dataset
Classifier train time (s) 313 Train/val split
Test acc. (%) 93.10% CPU-only
Table XIII summarizes the key timing and accuracy metrics. Feature extraction times are measured over the full CIFAR-10 dataset (45k train, 5k validation, 10k test), and classifier training times are measured on the train/validation split only. In terms of computation, the handcrafted SVM pipeline offers a competitive classical baseline by using compact 1,281-dimensional features with modest CPU-only costs (898 s feature extraction + 4,950 s training), clearly improving over pixel-based SVMs. Moreover, when fused with MobileNetV2 embeddings, a linear SVM reaches the best accuracy (93.10% vs. 92.43% for MobileNetV2 alone) while requiring only 313 s of classifier training compared to 23,054 s for end-to-end CNN fine-tuning, highlighting the time-efficiency of our hybrid approach.
As future work, we plan to explore more systematic hybrids of deep and handcrafted features by varying both the backbone (e.g., other CNNs and ViT variants) and the feature set (e.g., alternative entropy measures and color-space variants of HOG/LBP). It would also be interesting to study automated feature selection or compression on top of the proposed descriptor to further optimize the accuracy–efficiency trade-off and to validate the approach on additional, larger-scale benchmarks.
IX Conclusion
@@
To conclude, we presented a unified handcrafted feature extraction pipeline that integrates diverse descriptors, including temporal-complexity measures such as permutation entropy, spatial correlations, edge- and texture-based features (e.g., HOG, LBP), to construct a rich and interpretable feature representation. By applying this approach to a suite of benchmark datasets, including Fashion-MNIST, KMNIST, EMNIST, and CIFAR-10, we demonstrated that shallow classifiers, particularly SVM, can achieve competitive performance without relying on deep learning architectures.
Our results show that combining complementary handcrafted features significantly outperforms traditional single-feature methods in the classical machine learning regime (Fig. 6). Notably, we achieve up to 91.23% test accuracy on Fashion-MNIST and 64.81% on CIFAR-10 using only interpretable, resource-efficient techniques. Moreover, beyond serving as a strong classical baseline, the proposed handcrafted descriptors also complement existing deep learning architectures. For example, when fused with MobileNetV2 embeddings, a linear SVM achieves the best overall accuracy (93.10%) compared to MobileNetV2 alone (92.43%) on CIFAR-10. This improvement is obtained with substantially lower classifier training cost (313 s for the linear SVM) than end-to-end CNN fine-tuning (23,054 s), while preserving interpretability in the handcrafted component. Overall, these results reinforce that feature engineering remains relevant, not only as an efficient alternative in constrained settings, but also as a practical means to enhance deep models. This suggests hybrid strategies will remain valuable in future vision systems.
@@
Furthermore, in the future we plan to use these features as input representations for designing next-generation neural networks. Instead of feeding raw pixels into a deep model, we could build architectures that learn from more meaningful, structured inputs, like the features we extract here. This could lead to models that are faster to train, more accurate with less data, and easier to interpret. Therefore, combining the strengths of handcrafted features with the flexibility of deep learning could pave the way for next-generation computer vision systems that are both efficient and explainable.
Appendix A FashionMNIST and CIFAR-10 workflow (Algorithm box and parameters)
Algorithm 1: Fashion-MNIST workflow
   Input: $(X_{\mathrm{train}},y_{\mathrm{train}})$, $(X_{\mathrm{test}},y_{\mathrm{test}})$; PE $(m,\tau)$; patch $(p,s)$; diagonal offsets $\mathcal{K}$; HOG/LBP; SVM grid $\{C,\gamma\}$. Output: tuned RBF-SVM; test accuracy. 1. Define permutation entropy $PE(\mathbf{t};m,\tau)$ on a 1D sequence $\mathbf{t}$ (normalized by $\log_{2}(m!)$). 2. For each image $I$: (a) Row/column directional PE: $\sqrt{PE(\text{forward})\cdot PE(\text{reverse})}$ for all rows/columns. (b) Adjacent-row and adjacent-column Pearson correlations; set undefined values (NaN) to 0. (c) Diagonal and anti-diagonal PE over offsets $k\in\mathcal{K}$ using $\mathrm{diag}(I,k)$ and $\mathrm{diag}(\mathrm{fliplr}(I),k)$. (d) Patch-wise PE using a $p\times p$ window and stride $s$ on vectorized patches. (e) HOG descriptor and uniform-LBP normalized histogram. (f) Concatenate all blocks into a fixed-length vector $\phi(I)$. 3. Stack $\phi(I)$ into $\Phi_{\mathrm{train}}$ and $\Phi_{\mathrm{test}}$. 4. Fit StandardScaler on $\Phi_{\mathrm{train}}$ and transform train/test. 5. Tune $(C,\gamma)$ via CV over the predefined grid; train final SVM on the full standardized training set. 6. Evaluate accuracy on the standardized test set.
TABLE XIV: Parameters for the Fashion-MNIST pipeline
Component Setting (value)
Dataset Fashion-MNIST: train 60,000; test 10,000; grayscale $28\times 28$
Feature parallelism joblib Parallel: n_jobs = 50 (verbose progress)
Permutation Entropy (PE) order $m=3$, delay $\tau=1$; normalized by $\log_{2}(m!)$
Row/col PE aggregation $\sqrt{PE(\text{forward})\cdot PE(\text{reverse})}$ per row/column
Adjacent correlations Pearson between adjacent rows/cols; NaN $\rightarrow 0$
Diagonal PE offsets $\mathcal{K}=\{-10,-9,\dots,10\}$; mean of forward/reverse PE
Anti-diagonal PE same as diagonal after left-right flip of image
Patch-wise PE patch size $p=4$, stride $s=2$; PE on vectorized patch
HOG pixels_per_cell $(4,4)$; cells_per_block $(1,1)$; orientations = 9; feature_vector=True
LBP uniform LBP: $P=16$, $R=2$; histogram normalized
LBP histogram bins $P+2=18$ bins (implemented via bins $\{0,\dots,P+1\}$)
Feature dimensionality Entropy/corr block: $321$; HOG block: $441$; LBP block: $18$
Total per image: $780$
Standardization StandardScaler (fit on train, applied to train/test)
Classifier RBF-SVM (SVC); probability=True; random_state=42
Hyperparameter grid $C\in\{10,50,100,200\}$; $\gamma\in\{0.01,0.005,0.001\}$
Model selection HalvingGridSearchCV; CV folds = 3; scoring = accuracy
[…]
Search parallelism HalvingGridSearchCV: n_jobs = 50
Label noise Default 0.02, but reported run uses noise_level=0 (disabled)
Artifacts saved scaler and tuned SVM model (joblib dump)
Algorithm 2: CIFAR-10 RGB workflow
   Input: $(X_{\mathrm{train}},y_{\mathrm{train}})$, $(X_{\mathrm{test}},y_{\mathrm{test}})$; PE $(m,\tau)$; patch $(p,s)$; HOG/LBP settings; SVM grid $\{C,\gamma\}$. Output: tuned RBF-SVM; test accuracy. 1. Define permutation entropy $PE(\mathbf{t};m,\tau)$ on a 1D sequence $\mathbf{t}$ (normalized by $\log_{2}(m!)$). 2. For each RGB image $x$: (a) Split $x$ into channels $(x^{(R)},x^{(G)},x^{(B)})$. (b) For each channel image $I\in\{x^{(R)},x^{(G)},x^{(B)}\}$: i. Row/column directional PE: $\sqrt{PE(\text{forward})\cdot PE(\text{reverse})}$ for all rows/columns. ii. Adjacent-row and adjacent-column Pearson correlations; set undefined values (NaN) to 0. iii. Patch-wise PE using a $p\times p$ window and stride $s$ on vectorized patches. iv. Patch-wise PE on the horizontally flipped image $\mathrm{fliplr}(I)$. v. HOG descriptor and uniform-LBP normalized histogram. vi. Concatenate all blocks into a fixed-length channel vector $\phi(I)$. (c) Concatenate channel vectors to form the final RGB descriptor $\phi(x)=[\phi(x^{(R)}),\phi(x^{(G)}),\phi(x^{(B)})]$. 3. Stack $\phi(x)$ into $\Phi_{\mathrm{train}}$ and $\Phi_{\mathrm{test}}$. 4. Fit StandardScaler on $\Phi_{\mathrm{train}}$ and transform train/test. 5. Tune $(C,\gamma)$ via CV over the predefined grid; train final SVM on the full standardized training set. 6. Evaluate accuracy on the standardized test set.
TABLE XV: Parameters for the CIFAR-10 RGB pipeline
Component Setting (value)
Dataset CIFAR-10: train 50,000; test 10,000; RGB $32\times 32\times 3$
Feature parallelism joblib Parallel: n_jobs = 50 (verbose progress)
Permutation Entropy (PE) order $m=3$, delay $\tau=1$; normalized by $\log_{2}(m!)$
Row/col PE aggregation $\sqrt{PE(\text{forward})\cdot PE(\text{reverse})}$ per row/column
Adjacent correlations Pearson between adjacent rows/cols; NaN $\rightarrow 0$
Patch-wise PE patch size $p=8$, stride $s=4$; PE on vectorized patch
Flipped patch PE same patch-wise PE on $\mathrm{fliplr}(I)$
HOG (per channel) pixels_per_cell $(8,8)$; cells_per_block $(1,1)$; orientations = 9; feature_vector=True
LBP (per channel) uniform LBP: $P=8$, $R=1$; histogram normalized
LBP histogram bins $P+2=10$ bins (implemented via bins $\{0,\dots,P+1\}$)
Feature dimensionality Per channel: entropy/corr block $273$; HOG block: $144$; LBP block: $10$
Total per channel: $427$; Total per image: $1281$
Standardization StandardScaler (fit on train, applied to train/test)
Classifier RBF-SVM (SVC); probability=True; random_state=42
Hyperparameter grid $C\in\{10,50,100,200\}$; $\gamma\in\{0.01,0.005,0.001\}$
Model selection HalvingGridSearchCV; CV folds = 3; scoring = accuracy
[…]
Search parallelism HalvingGridSearchCV: n_jobs = 50
Label noise Default 0.02, but reported run uses noise_level=0 (disabled)
Artifacts saved scaler and tuned SVM model (joblib dump)